You have a planning horizon of H = 3 years and wish to immunize your investment for that horizon. You attempt to do so by buying a perpetual bond that pays $62 annually and has a YTM of 18% p.a. You will reinvest the coupons throughout this 3-year period and, additionally, you will sell the bond at the end of that 3-year period. Find the total cash flow you will have 3 years from now if interest rates remain unchanged throughout your planning horizon.
Answer #1

Future Value of annuity = A* ((1+rate)^n-1)/rate

Cash flow after 3 years = 62*((1+18%)^3-1)/18%

=62*3.5724

=$221.49